### v3

# MMIDR: Teaching Large Language Model to Interpret Multimodal Misinformation via Knowledge Distillation

Abstract.
Automatic detection of multimodal misinformation has gained a widespread attention recently. However, the potential of powerful Large Language Models (LLMs) for multimodal misinformation detection remains underexplored. Besides, how to teach LLMs to interpret multimodal misinformation in cost-effective and accessible way is still an open question. To address that, we propose MMIDR, a novel framework for multimodal misinformation interpretation and distillation reasoning, aiming to teach LLMs in providing fluent and high-quality textual explanations for their decision-making process. To convert multimodal misinformation into an appropriate instruction-following format, we present a data augmentation perspective and pipeline. This pipeline consists of a visual information processing module and an evidence retrieval module. Subsequently, we prompt the proprietary LLMs with processed contents to extract rationales for interpreting the authenticity of multimodal misinformation. Furthermore, we design an efficient knowledge distillation approach to distill the capability of proprietary LLMs in explaining multimodal misinformation into open-source LLMs. To explore several research questions regarding the performance of LLMs in multimodal misinformation detection tasks, we construct an instruction-following multimodal misinformation dataset and conduct comprehensive experiments. The experimental findings reveal that our MMIDR exhibits sufficient detection performance and possesses the capacity to provide compelling rationales to support its assessments.
multimodal misinformation detection; large language model; knowledge distillation; social media
Equal Contribution
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2024; XX, XXisbn: 978-1-4503-XXXX-X/24/06ccs: Information systems Multimedia information systemsccs: Information systems Social networks
1. Introduction
In the context of the widespread adoption of social media platforms and online news outlets, individuals enjoy the freedom to share daily information, express their opinions, and convey their emotions. However, this freedom accelerates the generation and spread of various forms of misinformation (e.g., fake news, rumors) (Zubiaga et al., 2018). Existing misinformation detection research has primarily focused on textual data (Yu et al., 2017; Nielsen and McConville, 2022). Nevertheless, most posts on these platforms are not confined to a specific modality. Detecting misinformation presented in different modalities poses a greater challenge, as it necessitates evaluating the credibility of each modality and their combinations (Abdali, 2022; Cao et al., 2020; Sharma et al., 2022).
In spite of notable advancements in comprehending and representing misinformation (Hu et al., 2022), existing models still face enormous challenges due to the intricate nature of the misinformation creation process. Misinformation creators demonstrate the ability to manipulate various aspects of information, employ diverse writing strategies, and operate with elusive underlying motives (Hu et al., 2023b). Leveraging evidence retrieved from the internet taps into a vast repository of world knowledge, enhancing the detection of meticulously crafted misinformation (Hu et al., 2023a). Notably, the retrieved evidence could be multimodal, encompassing both textual and visual elements, thereby posing a challenging task for conventional small language models (SLMs).
Recent advancements in the field of Large Language Models (LLMs), such as GPT-3 (Brown et al., 2020), InstructGPT (Ouyang et al., 2022), and GPT-4 (OpenAI et al., 2023), have demonstrated remarkable capabilities, such as the ability to follow instructions (Ouyang et al., 2022), perform knowledge-intensive tasks (Rubin et al., 2022; Shi et al., 2023), and address societal challenges(Jiang et al., 2023; Roy et al., 2023). In recent years, scholarly interest has been growing in investigating the application of LLMs in detecting misinformation (Buchholz, 2023; Li et al., 2023b; Pan et al., 2023; Bang et al., 2023; Caramancion, 2023; Chen and Shu, 2023). These studies aim to explore and develop effective methods that leverage the language understanding and reasoning abilities of LLMs to enhance the accuracy and robustness of misinformation detection. However, the potential of LLMs in generating explanations for misinformation remains underexplored. Previous works on the explainability of misinformation detectors have primarily relied on extraction methods (Shu et al., 2019; Lu and Li, 2020; Li et al., 2021; Yang et al., 2022). Nevertheless, LLMs exhibit the ability to generate fluent, natural language-based explanations for the given misinformation while predicting its authenticity. This capability is more human-friendly compared to earlier approaches (Chen and Shu, 2023). Therefore, how to teach LLMs to interpret misinformation remains an open question.
It has been demonstrated that LLMs can be naturally extended to multimodalities, such as GPT-4 (OpenAI et al., 2023), Gemini (Team et al., 2023), and GPT-4V (OpenAI, 2023). Multimodal Large Language Models (MLLMs) possess the ability to understand and process information across multiple modalities, surpassing the capabilities of traditional language-only LLMs. This broadened comprehension of information across multiple modalities enables MLLMs to closely simulate human perception, thereby expanding the potential for a diverse array of real-world applications. While proprietary LLMs and MLLMs exhibit remarkable capabilities, they still have certain limitations. One notable drawback is their limited accessibility and higher cost. In contrast, open-source LLMs, exemplified by LLaMA (Touvron et al., 2023), offer distinct advantages. Additionally, recent studies have widely adopted knowledge distillation techniques, confirming their effectiveness in bridging the performance gap between proprietary and open-source LLMs (Hsieh et al., 2023; Yang et al., 2023; Jung et al., 2023; Ramnath et al., 2023; Liu et al., 2023; Xu et al., 2024). To the best of our knowledge, the research on utilizing LLMs and MLLMs for multimodal misinformation detection via knowledge distillation techniques remains underexplored.
Taking the considerations above, we propose MMIDR, a novel framework for Multimodal Misinformation Interpretation and Disti- llation Reasoning. We present a data augmentation perspective and pipeline to convert image-text pairs into an appropriate instruction-following format. Specifically, for a given multimodal misinformation, we utilize OCR and image captioning techniques to process visual information. Additionally, we obtain both visual and textual evidence through evidence retrieval (Hu et al., 2023a). After that, we feed the processed contents into proprietary LLMs (e.g., ChatGPT (Ouyang et al., 2022)) and extract relevant rationales aimed at explaining the authenticity of multimodal misinformation. Finally, we design an efficient knowledge distillation approach. By integrating the original multimodal information content with the rationales provided by proprietary LLMs, also known as teacher LLMs, we employed the LoRA (Low-Rank Adaptation) (Hu et al., 2021) technology to fine-tune open-source LLMs (e.g., LLaMA (Touvron et al., 2023), MiniGPT-v2 (Chen et al., 2023)), also known as student LLMs. In this way, MMIDR can distill the capability of teacher LLMs in explaining multimodal misinformation into student LLMs. This process can promote a comprehensive understanding and integration of multimodal misinformation content by open-source LLMs, empowering them to generate high-quality rationales to interpret their decision-making processes.
The main contributions of this paper are as follows:
  • We investigate the utilization of LLMs and MLLMs for the detection of multimodal misinformation. To our best knowledge, this is the first exploration in this area.
  • We propose MMIDR, a framework designed to teach LLMs to interpret multimodal misinformation, aiming to provide fluent and high-quality textual explanations for their decision-making process.
  • We introduce a data augmentation perspective and pipeline that converts multimodal retrieval-enhanced misinformation data into an appropriate instruction-following format. This process enables the acquisition of knowledge from proprietary LLMs, leading to the construction of a dataset $\text{MR}_{llm}^{2}$.
  • We propose an efficient knowledge distillation approach designed to address the limited accessibility and higher cost of proprietary LLMs. Simultaneously, our approach aims to bridge the performance gap between open-source LLMs and proprietary LLMs, empowering the former to generate high-quality explanations.
  • We conduct extensive experiments on the dataset $\text{MR}_{llm}^{2}$, investigating several research questions regarding the performance of LLMs in multimodal misinformation detection tasks. Experimental results demonstrate that our framework has sufficient detection performance and can provide compelling rationales to support its assessments.
2. Related Works
2.1. Multimodal Misinformation Detection
Existing misinformation Detection methods (Qian et al., 2018; Bhattarai et al., 2021; Guo et al., 2018; Potthast et al., 2017; Conroy et al., 2015)mostly fall into text-based. Some previous works have explored the explainability of misinformation detectors. dEFEND (Shu et al., 2019) leverages a combined analysis of news content and user comments to capture explainable user comments for misinformation detection. GCAN (Lu and Li, 2020) employs a dual co-attention mechanism to better learn the representations of user interactions, retweet propagation, and their correlation, thereby enabling the generation of reasonable explanations. However, previous works on the explainability of misinformation detectors have primarily relied on extraction methods, which may lack human-friendliness. In this paper, we effectively leverage LLMs to generate fluent, natural language-based explanations for interpreting misinformation
More recently, several methods (Wang et al., 2018; Khattar et al., 2019; Wu et al., 2021; Chen et al., 2022; Singhal et al., 2022; Sun et al., 2023; Wang et al., 2023) have been proposed to evaluate the credibility of each modality and their combinations, achieving superior performance in multimodal misinformation detection. For instance, EANN (Wang et al., 2018) employs an adversarial approach to remove event-specific features, thereby enhancing the adaptability to newly emerging events. KDCN (Sun et al., 2023) suggests a unified framework for detecting multimodal misinformation by modeling cross-modal and content-knowledge inconsistencies. CAFE (Chen et al., 2022) measures cross-modal ambiguity through the assessment of Kullback-Leibler (KL) divergence between unimodal feature distributions. COOLANT (Wang et al., 2023) leverages cross-modal contrastive learning to achieve more accurate image-text alignment. However, these approaches focus solely on the multimodal information without exploiting the internet to retrieve evidence, which could enhance detection performance (Hu et al., 2023a). Given that the retrieved evidence may consist of both textual and visual components, conventional small language models face challenges in processing such multimodal data. In this study, we advocate the utilization of large language models to tackle these intricate scenarios.
2.2. LLMs for Detecting Misinformation
In recent years, there have been increasing efforts to investigate the application of LLMs in detecting misinformation (Buchholz, 2023; Li et al., 2023b; Pan et al., 2023; Bang et al., 2023; Caramancion, 2023; Chen and Shu, 2023; Pelrine et al., 2023; Pavlyshenko, 2023; Cheung and Lam, 2023; Pan et al., 2023; Zhang and Gao, 2023). These studies can be divided into three methods: 1) directly prompting-based methods, 2) fine-tuning-based methods, and 3) augmentation-based methods, which augment LLMs with external knowledge or tools. For instance, ARG (Hu et al., 2023b) employs a small model to encode specific news along with the textual and common-sense descriptions generated by LLMs. DELL (Wan et al., 2024) utilizes LLMs to generate synthesized reactions and comments from diverse viewpoints, employing them for proxy tasks related to prediction and explanation. (Pavlyshenko, 2023) utilize LoRA to fine-tune LLaMA2 (Touvron et al., 2023) aiming to adapt it for various tasks. FACTOOL (Chern et al., 2023) proposes a multi-task and multi-scenario framework aimed at detecting factual errors in LLM-generated texts. While LLMs can naturally expand to multimodalities, the extensive potential they hold for combating multimodal misinformation in real-world settings remains largely unexplored.
2.3. Instruction Tuning
Recent studies (Ouyang et al., 2022; Wang et al., 2022a, b) have explored methods for refining the instruction-following capabilities of LLMs. These studies reveal that by fine-tuning on specific instruction-following datasets, LLMs can follow natural language instructions, enabling them to perform various real-world tasks effectively. More importantly, there has been a surge in the adoption of knowledge distillation techniques (Hsieh et al., 2023; Yang et al., 2023; Jung et al., 2023; Ramnath et al., 2023; Liu et al., 2023; Gou et al., 2021; Gupta and Agrawal, 2022; Xu et al., 2024), focusing on distilling off-the-shelf language models using instructional data obtained from strong LLMs. This trend aims to enhance the zero- and few-shot abilities of language models. For instance, the Evol-Instruct method (Xu et al., 2023) expands the instructions in two dimensions: difficulty, such as the rephrasing of questions to heighten complexity, and diversity, entailing the generation of long-tailed instructions. This method has been used to expand the distillation of coding (Luo et al., 2023b) and math (Luo et al., 2023a). In the vision-language domain, LLaVA (Liu et al., 2024) translates images into textual descriptions, including captions and bounding boxes, and subsequently employs GPT-3.5 and GPT-4 for distillation to generate instruction-following data related to visual content. In this paper, we utilize GPT-3.5 to extract rationales and construct an instruction-following multimodal misinformation dataset. We employ LoRA to train student LLMs (e.g., LLaMA, MiniGPT-v2), enabling them to interpret multimodal misinformation effectively.
Figure 1. Model Architecture Overview of MMIDR. Given multimodal information, we first convert it into an appropriate instruction-following format. Then we employ the labeling template to prompt the teacher LLM in order to obtain rationales for the authenticity labels assigned to provided multimodal information. Finally, we employ LoRA (Hu et al., 2021) to train the student LLM on the integration of the original multimodal information and the rationales. $I(I+V)$ denotes that the input for student LLM could be textual instruction ($I$) or both textual instruction and image ($I+V$) based on whether student is multimodal.
3. Methodology
In this section, we introduce our novel framework MMIDR, designed specifically to teach LLMs to interpret multimodal misinformation. The primary objective of MMIDR is to provide fluent and high-quality textual explanations for the decision-making process. The overall model structure is illustrated in Figure 1. For each given multimodal information, we first conduct data augmentation (§3.1), which involves transforming the multimodal retrieval-enhanced misinformation into an appropriate instruction-following format. Subsequently, we input this processed content into teacher LLMs for the extraction of rationales (§3.2). Finally, we efficiently perform knowledge distillation to distill the capability of teacher LLMs in explaining multimodal misinformation into student LLMs (§3.3).
3.1. Data Augmentation
For each input multimodal information $\mathbf{P}=\{T,V\}\in\mathcal{D}$, where $T$, $V$, and $\mathcal{D}$ represent text, image, and the original dataset, respectively. This work requires the application of large language models to guide and extract rationales (§3.2). To facilitate this, we convert the original multimodal information’s image $V$ into textual information compatible with the large language model (§3.1.1). Subsequently, we use text $T$ and image $V$ separately for retrieval purposes, aimed to procure evidence that supports the large language model’s reasoning process (§3.1.2).
3.1.1. Visual Information Processing
Given visual content $V$, we leverage Optical Character Recognition (OCR) technology to discern and transmute characters embedded within the image into an editable text format, designated as $V_{ocr}$. This process is specifically executed through the utilization of Baidu’s PaddleOCR framework [^1] https://github.com/PaddlePaddle/PaddleOCR.Simultaneously, we employ Image Captioning technology to provide descriptions for the image content. In particular, we choose to employ the BLIP-2 $\text{OPT}_{2.7B}$ model(Li et al., 2023a) for generating descriptive text, resulting in $V_{caption}$.
3.1.2. Evidence Retrieval
Through the retrieval of evidence from the internet, the model is empowered to embrace a broader scope of global knowledge, thereby facilitating the identification of meticulously crafted misinformation. Drawing inspiration from the evidence retrieval methodology(Hu et al., 2023a), our approach involves utilizing the image $V$ from the input post as a query. By employing the Google Reverse Image [^2] https://cloud.google.com/vision, we systematically retrieve textual evidence denoted as $E_{T}$. The search engine yields a selection of images resembling the query image, from which we extract textual evidence, including titles and descriptions.
[…]
3.2. Rationales Elicitation
Rationales elicitation refers to using a teacher LLM to generate explanations for the given multimodal information. Specifically, each instance is represented as a tuple $x=\{T,V_{ocr},V_{caption},E^{m}_{T},E^{n}_{V}\}$, comprising the textual content of the post $T$, OCR-recognized text derived from visual content $V_{ocr}$, text generated from image captions associated with the visual content $V_{caption}$, along with a collection of textual evidence $E^{m}_{T}$ and visual evidence $E^{n}_{V}$, where $m$ and $n$ denote the number of pieces of evidence.
To utilize the instruction-following ability of LLMs (Wei et al., 2021), we design a simple labeling prompt template $I$ that prompts the teacher LLM to generate rationales:
Now give you a multimodal post and the corresponding textual evidences and image evidences, these evidences may be relevant or irrelevant. Each post is a tuple of (text, image_ocr, image_caption). Evidences are a list of sentences, which are split by ”¡and¿”. ## Three-way classification scheme Each post is labeled in three-way classification scheme: [0: ”non-rumor”, 1: ”rumor”, 2: ”unverified”]. ## Fine-grained labels ¡Non-rumor — Rumor — Unverified fine-grained labels¿ ## Post Post:[ text: {post_text}, image ocr: {image_ocr}, image caption: {image_caption}] ## Evidences Textual_evidence: {textual_evidence} Image_evidence: {visual_evidence} ## Label label: ¡Non-rumor — Rumor — Unverified¿ ## Explanation Now, you are required to explain why the post is labeled as ¡Non-rumor — Rumor — Unverified¿ according to the evidences, and decide which fine-grained labels are suitable for the post.
To instantiate the template, each instance $x$ is incorporated into the designated position within the template’s blue segment. Simultaneously, based on its ground-truth label $g$, the template’s green segment is specifically populated. The comprehensive details regarding fine-grained labels (Hu et al., 2023a) can be found in Table 1. An instantiated template is denoted as $I(x,g)$.
By feeding the instantiated template $I(x,g)$ into the teacher LLM, we could annotate the provided input $x$ through the labeling prompt template template, yielding the output $y$ of the rationale. This process can be formulated as follows:
(1) $\mathcal{D}^{\text{(rationales)}}=\{x,y|x\sim\mathcal{D},y\sim p_{T}\left(y|I(x,g)\right)\}.$
where $I(x,g)$ is a template instantiated by the given instance $x$ and the ground-truth label $g$. $p_{T}$ represents the teacher LLM with parameters $\theta_{T}$. Additionally, to guarantee the output is ended with the prediction of the label, each output $y$ is appended with a specific text of ”Therefore, the post is labeled as ¡label¿”. Therefore, the output $y$ encompasses both the rationale and the ground-truth label, which could be used to train the student LLM subsequently.
Table 1. Mappings from fine-grained labels to three standardized labels.
Standardized Fine-grained Labels
Rumor Altered, Altered Image, Altered Photo, Bad Math, Bad Science, Clickbait, Commontion, Death Hoax, Distorts the Facts, Doctored,
FAKE, Fake News, Fake Quotes, Fake Quoto, Fake Tweet, False, False and Misleading, False Headline, False!, Flawed-Reasoning, Flip-flop,
Full Flop, Hoax, Hoax!, Inaccurate, Incorrect, It’s A Joke, Lacks Context, Likely False, Misattributed, Misleading, Misleading and False,
Misleading!, Misleading., Misplaced Context, Misrepresented, Missing Context, Mixed, Mixture, Mostly False, Needs Context, Not Legit,
Not the full story, Not the Whole Story, Out of Context, Pants on Fire, Partly False, Photoshopped, Selective, Spins the Facts, Staged Skit,
Suspicious, This claim is False., This is misleading., This is not true., Three Pinocchios, Totally Fake, Totally False, Trolling,
Two Pinocchios, Wrong, Wrong Number, Wrong Numbers
Non-Rumor Accurate, Close to the mark, Correct, Mostly correct, Mostly True, Mostly-Accurate, No Fraud, Partially True, Partially-Correct, True
Unverified Baseless, Myth, No Evidence, No Proof, Not Supported, This lacks evidence., Unclear, Unproven, Unsubstantiated, Unsupported
3.3. Knowledge Distillation
Knowledge distillation refers to fine-tune the student LLM by maximizing the likelihood of sequences generated by the teacher LLM, ensuring alignment between the predictions of the student and the teacher. Based on the dataset $\mathcal{D}^{\text{(rationales)}}$, we tune the off-the-shelf language model, adhering to its original auto-regressive training objective. To facilitate this tuning process, we employ LoRA (Hu et al., 2021) optimization, wherein the language model is frozen, and optimization is solely applied to the rank decomposition components of the Transformer layers.
Due to the absence of label during inference, we design a new instruction template $I^{\prime}$ derived from the labeling prompt template $I$ by removing label-related contents, i.e. the green segments in the template. That is, instantiating the $I^{\prime}$ only requires the instance $x$, i.e. $I^{\prime}(x)$. For an output $y$ comprising $L$ tokens, the learning objective of student LLM is to maximize the conditional language modeling:
(2) $p\left(y\mid I^{\prime}(x)\right)=\prod_{i=1}^{L}p_{\theta}\left(y_{i}\mid I^{\prime}(x),y_{1:i-1}\right)$
where $I^{\prime}(x)$ denotes the instantiated instruction, and $\theta$ represents the trainable parameters. Considering that the student LLM adopted in our study encompasses multimodal large language models, its input component is equipped to accommodate the original visual content. Therefore, the aforementioned process can be described as follows:
(3) $p\left(y\mid V,I^{\prime}(x)\right)=\prod_{i=1}^{L}p_{\theta}\left(y_{i}\mid V,I^{\prime}(x),y_{1:i-1}\right)$
where $V$ is the original visual content.
4. Experiments
In our experimental endeavors to explore the utilization of LLMs in the realm of multimodal misinformation detection, we aim to address the following research questions:
  • RQ1: Do closed-source LLMs and open-source LLMs possess sufficient background knowledge to detect multimodal misinformation?
  • RQ2: Can fine-tuned open-source LLMs independently serve as multimodal misinformation detectors?
  • RQ3: After knowledge distillation, do open-source LLMs demonstrate sufficient detection performance and provide compelling rationales to support their assessments?
4.1. Experimental Configurations
4.1.1. Datasets
The experimental dataset is constructed based on the original dataset $\text{MR}^{2}$ (Hu et al., 2023a). The dataset comprises two modalities, encompassing both images and text, and incorporates two languages, English and Chinese. By leveraging the remarkable cross-lingual capacities inherent in LLMs, the dataset developed in our study no longer strictly segregates between languages.
[…]
Table 2. Statistics of $\text{MR}_{llm}^{2}$ dataset and the labels distribution.
Set Non_rumor Rumor Unverified Total
Train 3,592 2,392 5,200 11,184
Test 421 280 608 1,309
Total 4,013 2,672 5,808 12,493
Figure 2. The length distribution of training data composed of different numbers of evidence.
4.1.2. Evaluation Metrics and Baselines
The evaluation metrics include Accuracy, Precision, Recall, and F1-score. We compare MMIDR with 3 baseline model for multimodal misinformation detectio: 1) BERT (Devlin et al., 2018)+ ResNet (He et al., 2016), 2) CAFE (Chen et al., 2022), leveraging cross-modal ambiguity measurement for adaptive aggregation, and 3) COOLANT (Wang et al., 2023), employing cross-modal contrastive learning to enhance image-text alignment accuracy.
4.1.3. Implementation Details
For MMIDR, we use LLaMA2-Chat-7B (Touvron et al., 2023) as the LLM backbone, which was pretrained on 2 trillion tokens of data from publicly available sources. We employ MiniGPT-v2 (Chen et al., 2023) as the foundational architecture for our MLLM. MiniGPT-v2 directly assimilates visual tokens from a ViT vision encoder (Fang et al., 2023) and projects them into the feature space of an LLM (Touvron et al., 2023).
Based on the constructed data, we perform fine-tuning of language models employing LoRA technology (Hu et al., 2021). Specifically, we equip the projection layers of query, value, and output with LoRA layers. The LoRA attention dimension and scaling alpha are set to 16. While maintaining the language model in a frozen state, the LoRA layers are optimized using the AdamW (Loshchilov and Hutter, 2018). All models are fine-tuned over 50 epochs, with a batch size 64. The learning rate is set to $1\times 10^{-4}$, and the maximum length of newly generated tokens is constrained to 2048.
4.2. RQ1: Zero-shot Experiments
In this section, we perform zero-shot experiments to verify whether large language models inherently possess enough background knowledge to detect multimodal misinformation. Specifically, we evaluate both ChatGPT (gpt-3.5-turbo) (Ouyang et al., 2022) and MiniGPT-v2 (Chen et al., 2023) on the test set of $\text{MR}_{llm}^{2}$. The results are presented in Table 3.
From the results, we can notice that the performance of all LLMs does not attain a satisfactory level [^4] Detailed comparisons are not conducted here due to the limited compatibility of LLaMA2-7B-Chat (Touvron et al., 2023) with instructions.. With evidence support, the accuracy of the teacher model gpt-3.5-turbo is below 50%. The classification performance of MiniGPT-v2 is slightly better than that of gpt-3.5-turbo, suggesting that the introduction of original visual content contributes to the detection of multimodal misinformation. It is evident that even powerful LLMs may lack sufficient background knowledge to detect multimodal misinformation.
Table 3. Zero-shot Experiments with gpt-3.5-turbo and MiniGPT-v2.
Models Accuracy Precision Recall F1
gpt-3.5-turbo 48.89 42.91 42.04 41.56
- w/o textual evidence 49.89 42.13 42.25 41.18
- w/o visual evidence 49.73 42.98 40.42 38.30
- w/o any evidence 36.75 38.54 32.48 32.94
MiniGPT-v2 53.72 53.70 53.38 52.52
Additionally, it is noted that furnishing either solely textual evidence or visual evidence yields a marginally higher accuracy for the teacher model, compared to the simultaneous presentation of all evidence. This phenomenon may be attributed to the potential challenge posed to the teacher model in processing an excessive volume of evidence. Therefore, we initiated a comprehensive examination to discern the influence of evidence quantity on the classification accuracy of the teacher model. In Figure 3, we vary the numbers of retrieved visual and textual evidence from $1\sim 10$ and report the Accuracy on the test set. It is evident that the quantity of evidence significantly impacts the decisions made by LLMs, reaching an optimal effect when the cumulative quantity of visual and textual evidence is 6, as depicted by the red and blue lines. Nevertheless, these effects do not surpass the performance achieved by MiniGPT-v2 (53.72%). Furthermore, the fluctuation results further indicate the challenges faced by LLMs in effectively detecting multimodal misinformation.
Figure 3. The classification accuracy of the teacher model using different numbers of evidence.
4.3. RQ2: Performance Comparison and Ablation Study
In this section, we compare MMIDR with other baselines on multimodal misinformation detection tasks. Note that $\text{MMIDR}_{\text{LLaMA2}}$ and $\text{MMIDR}_{\text{MiniGPT-v2}}$ denote the results obtained by fine-tuning MMIDR model on LLaMA2 (Touvron et al., 2023) and MiniGPT-v2 (Chen et al., 2023), respectively. We also compare two variants of MMIDR: 1) w/o evidence&rationale, where we remove evidence from the instruction during the detection of multimodal misinformation. Given that rationales encompass descriptions of certain evidence, they should be excluded in this context. 2) w/o rationale, where we simply eliminate the rationales. Hence, these two variants can be considered equivalent to inputting multimodal misinformation and having the LLMs directly generate the veracity label. The results are shown in Table 4. From the results, we observe that:
  • The fine-tuned baseline models show a significant improvement in accuracy compared to the zero-shot experiments of LLMs, demonstrating that smaller models can achieve higher classification performance after task-specific training.
  • After being fine-tuned with MMIDR w/o rationale, both LLaM-A2 and MiniGPT-v2 achieve optimal performance, reaching 95.83% and 95.95%, respectively. Compared to the baseline model BERT+ResNet, there is an improvement of over 20%. This demonstrates that LLMs have the capability to handle multimodal misinformation detection tasks. Furthermore, the inherent robust background knowledge of LLMs enables them to adapt effectively to various tasks with minimal information.
  • After being fine-tuned with MMIDR w/o evidence&rationale, the performance is slightly worse compared to MMIDR w/o rationale (95.57% vs. 95.83% & 94.19% vs. 95.95%), further confirming that evidence has a certain effect on enhancing the model’s detection performance.
  • In the main experimental section of MMIDR, there is a slight performance decline compared to the two aforementioned variants. We attribute this to two factors: 1) The two aforementioned variants exclusively employ LLMs for a relatively straightforward three-classification task; and 2) The objective function during the distillation process primarily emphasizes the alignment of predictions between the student model and the teacher model, omitting the inclusion of task-specific classification loss. Nevertheless, our primary goal is the distillation of open-source LLMs/MLLMs through the utilization of rationales generated by the teacher model, facilitating the production of high-quality, cohesive textual explanations for their decision rationale. In pursuit of this objective, a marginal performance decrease is deemed acceptable (93.63% vs. 95.83% & 94.04% vs. 95.95%).
Table 4. Performance comparison between MMIDR and other methods on $\text{MR}_{llm}^{2}$ dataset. The best two results are bolded and underlined, respectively. For the main experimental results of the MMIDR, apply red font for annotation.
Models Accuracy Precision Recall F1
gpt-3.5-turbo(zero-shot) 48.89 42.91 42.04 41.56
MiniGPT-v2(zero-shot) 53.72 53.70 53.38 52.52
BERT+ResNet 75.40 75.34 72.51 73.61
CAFE 78.69 78.28 74.39 75.71
COOLANT 82.97 84.88 77.49 79.56
$\text{MMIDR}_{\text{LLaMA2}}$ 93.63 93.62 93.62 93.62
- w/o evidence&rationale 95.57 95.50 95.66 95.56
- w/o rationale 95.83 95.80 95.87 95.83
$\text{MMIDR}_{\text{MiniGPT-v2}}$ 94.04 94.12 93.90 94.00
- w/o evidence&rationale 94.19 94.22 94.11 94.16
- w/o rationale 95.95 95.97 95.89 95.92
Table 5. Examples from $\text{MR}_{llm}^{2}$ (Chinese is translated into English). We annotate instructions for the LLMs, highlighting data augmentation content in blue, which includes OCR, captions, and evidence. Key evidence is indicated in brown. In the generated explanations, we mark in purple the parts that most clearly indicate misinformation.
Examples from $\text{MR}_{llm}^{2}$:
Input Post
天津一女子在路上扶起了摔倒的老人帮忙送去救治并垫付3000元,老人清醒后指认是女子撞到了他,老人的子女要求女子再拿10000元。女子报了警… (A woman in Tianjin helped an elderly person who had fallen on the road, took him for medical treatment, and paid 3,000 yuan. After the elderly person regained consciousness, he accused the woman of causing the accident. The elderly person’s children demanded an additional 10,000 yuan from the woman. The woman reported the incident to the police.)
While India is in denial about border tensions in Ladakh, China linked accounts shared this photo with the caption India China border, in which Indian soldiers are seen lying on the ground with hands and legs tied up with rope.
Instruction
Now give you a multimodal post and the corresponding textual evidences… image ocr: 老人跌倒路边一直求救女子暖心扶起帮老人找家我拉你一下好不好@历史课课代表, (The elderly person fell by the roadside, continuously calling for help. The compassionate woman helped the elderly person up and assisted in finding his way home, saying, ”Let me give you a hand, is that okay?” @HistoryClassRep) image caption:an old man lying on the ground with his head down… Evidences… ¡and¿ 雷音振长空01-0220:57 微博轻享版任何值得称道的正义之光,都应建立在事实之上。可惜的是,这个被这么多大v转发的新闻是假的 (Any commendable light of justice should be built upon facts. Unfortunately, the news widely shared by many influencers is fake.) … Now,you are required to explain…
Now give you a multimodal post and the corresponding textual evidences… image ocr: @Lz1277 87829, image caption: soldiers lay on the ground while others stand around them. Evidences… ¡and¿ 72 Indian soldiers injured, captured in clash with Chinese Army… ¡and¿ Footage claims to show ’Indian troops beating a captured Chinese soldier’ amid border tension… ## Explanation: Now, you are required to explain and label the post in three-way classification scheme: [”non-rumor”, ”rumor”, ”unverified”].
gpt-3.5-turbo
The post is labeled as rumor because the textual evidences show conflicting information and the image caption does not directly support the text. The textual evidence includes multiple instances of the post being shared on social media, but also includes a comment suggesting that the news is fake. Additionally, the image OCR does not provide any relevant information to confirm the story in the text. Fine-grained labels that are suitable for the post include ”Misleading and False” and ”Fake News” as the story is being disputed and there is a suggestion that it is not true. Therefore, the post is labeled as ¡label¿ rumor ¡/label¿.
The post is labeled as a rumor because the textual evidence mentions that the photo was circulated among Chinese military websites, and there are conflicting reports about the incident, with some claiming that Indian soldiers were captured and others claiming that Indian troops beat a captured Chinese soldier. This indicates that there is uncertainty and conflicting information surrounding the photo, leading to its classification as a rumor. As for the fine-grained labels, ”Altered Image” and ”Misleading” would be suitable for the post, as the photo is being shared with a misleading caption and there are conflicting reports about the incident depicted in the image. Therefore, the post is labeled as ¡label¿ rumor ¡/label¿.
$\text{MMIDR}_{\text{LLaMA2}}$
The post is labeled as rumor because the textual and image evidences do not provide clear and credible information about the incident. The textual evidence includes repeated mentions of ”Sina Visitor System,” which is not relevant to the content of the post. The image evidence includes a screenshot of a text message in Chinese, which does not provide any relevant information about the incident. Additionally, the image caption describes an old man lying on the ground, which is not directly related to the claim made in the post. Fine-grained labels that are suitable for the post include ”Misleading,” ”Lacks Context,” and ”Unverified,” as the post lacks clear and credible information, and the image evidence does not support the claim made in the text. Therefore, the post is labeled as ¡label¿ rumor ¡/label¿.
The post is labeled as a rumor because the textual evidence suggests that the photo of Indian soldiers being tied up with rope is being circulated among Chinese military websites, but there is no concrete evidence to support this claim. Additionally, the image OCR and caption do not provide any relevant information to verify the claim made in the text. Fine-grained labels that are suitable for the post include: Altered Image, Misleading, Misattributed, and False. Therefore, the post is labeled as ¡label¿ rumor ¡/label¿.
$\text{MMIDR}_{\text{MiniGPT-v2}}$
The post is labeled as a rumor because the textual evidence includes multiple screenshots of Chinese social media posts and text messages, but there is no concrete evidence or credible source to support the claim that the woman in the post helped an old man and paid 3000 yuan for his medical expenses. The image caption of an old man lying on the ground with his head down also does not provide any concrete evidence to support the claim. Fine-grained labels that are suitable for the post include ”Misleading”, ”Lacks Context”, and ”Unverified”. These labels reflect the lack of concrete evidence and the potential for misleading information in the post. Therefore, the post is labeled as ¡label¿ rumor ¡/label¿.
The post is labeled as rumor because the textual and image evidences suggest that there is a claim of Indian soldiers being tied up and beaten by Chinese soldiers, which is not verified. The textual evidence mentions a photo circulating among Chinese military websites showing PLA troops bringing down several Indian soldiers at a close-quarters fight at Pangong Lake border area with India. The image evidence does not directly support the claim, but it does not refute it either. Fine-grained labels that are suitable for the post include ”Misleading”, ”Misrepresented”, and ”Unverified”. These labels reflect the lack of verification and the potential for misleading information in the post. Therefore, the post is labeled as ¡label¿ rumor ¡/label¿
4.4. RQ3: Faithful Detection and Interpretable Results
To evaluate the ability of LLMs in producing high-quality and coherent text explanations for multimodal misinformation, we conducted a case study, as detailed in Table 5. We specifically chose one Chinese example and one English example for in-depth analysis.
  • Chinese Example (left): In this case, the image depicts an elderly man lying on the road. The accompanying text narrates the story of a woman who, driven by goodwill, comes to the aid of the elderly individual. Unfortunately, her good intentions lead to false accusations, compelling her to seek legal intervention by reporting the incident to the police. The most notable evidence includes a comment stating, ”Unfortunately, the news widely shared by many influencers is fake”.
    It is noteworthy that gpt-3.5-turbo successfully captured the crucial evidence mentioned above, pointing out ”a comment suggesting that the news is fake”. This demonstrates its exceptional expressive capabilities and comprehensiveness in serving as a teaching model. In contrast, $\text{MMIDR}_{\text{LLaMA2}}$ simply highlights the insufficiency of evidence, acknowledging that the image solely depicts an elderly individual on the ground, without corresponding to the textual description of a female figure. $\text{MMIDR}_{\text{MiniGPT-v2}}$ predominantly raises concerns about the payment of 3000 yuan by the female, concurrently suggesting that the image merely depicts the elderly person on the ground, failing to substantiate the events adequately described in the textual account. This finding indicates that, despite the fact that the distilled student model can generate fluent expressions and provide explanations to some extent, its ability to capture crucial information remains inferior to that of the teacher model.
  • English Example (right): In this case, the image depicts a group of soldiers lying on the ground. The textual context specifies the location as the India-China border, accompanying an image featuring Indian soldiers lying on the ground, with their hands and legs tied with ropes. The most significant evidence is someone claiming that these are Indian soldiers who have been captured, while others assert that it is the Indian army beating a captured Chinese soldier, presenting conflicting narratives.
    The crucial contradiction mentioned above is successfully captured by gpt-3.5-turbo, further demonstrating its ability to comprehend information. In comparison, $\text{MMIDR}_{\text{LLaMA2}}$ still deems there is insufficient evidence. $\text{MMIDR}_{\text{MiniGPT-v2}}$ asserts a general lack of verification and the potential presence of misleading information. However, it specifically references certain textual evidence related to the India-China conflict, which exhibits limited relevance to the content of the present post. This observation further highlights the remaining gap in explanatory ability of the distilled student model, providing a potential avenue for exploration in future research.
5. Conclusion
In this paper, we propose MMIDR, a framework designed to teach LLMs to provide fluent and high-quality textual explanations for their assessments. Given multimodal information, we first conduct data augmentation to convert the multimodal retrieval-enhanced misinformation into an appropriate instruction-following format. Subsequently, we feed the processed content into ChatGPT for the extraction of rationales. Finally, we design an efficient knowledge distillation approach, which employs LoRA to train student LLMs (e.g., LLaMA, MiniGPT-v2) on our constructed instruction-following multimodal misinformation dataset. Experimental results demonstrate that our MMIDR exhibits sufficient detection performance and possesses the capacity to generate compelling rationales to interpret its decision-making processes. Nevertheless, the distilled student model still falls short in capturing crucial information compared to the teacher model. We hope our work can facilitate the effective utilization of LLMs and MLLMs in detecting multimodal misinformation.